OpenStudy (anonymous):

Help? How do I solve this? x^2 + 11x + 18 = 5/6

OpenStudy (goten77):

what do you want to solve for? x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can factor out the x^2

OpenStudy (anonymous):

(x + 9)(x + 2) = 5/6

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now what do I do?

OpenStudy (alexwee123):

no you would subtract the 5/6 on both sides to make the whole thing=0 if you want to factor it

OpenStudy (anonymous):

I am trying to solve for x...

OpenStudy (anonymous):

w8 we cant factor out

OpenStudy (anonymous):

the original problem is

OpenStudy (alexwee123):

OH that really changes the problem :(

OpenStudy (anonymous):

I just really need help solving it.

OpenStudy (anonymous):

rather we can do this 108-5=103 x^2+11x+103/6

OpenStudy (alexwee123):